A company offers a starting salary of 35,000 with raises of $2,500 per year.Find the total salary over a ten-year period.
Answer by checkley79(3341)

35,000 THE FIRST YEAR
35,000+2,500=37,500 SECOND YEAR
37,500+2,500=40,000 THIRD YEAR
40,000+2,500=42,500 FORTH YEAR
42,500+2,500=45,000 FIFTH YEAR
45,000+2,500=47,500 SIXTH YEAR
47,500+2,500=50,000 SEVENTH YEAR
50,000+2,500=52,500 EIGHTH YEAR
52,500+2.500=55,000 NINTH YEAR
55,000+2,500=57,500 TENTH YEAR.